# ThinkVideo: High-Quality Reasoning Video Segmentation with Chain of Thoughts

### v2

@@ 3.2 ThinkVideo: Reasoning Video Instance Segmentation
Then, we use the reasoning image segmentation model $\mathcal{F}_{seg}$ to generate per-instance key masks $\{\tilde{m}_{i}\in\mathbb{R}^{H\times W}\}$ based on the the instance-level selectivity. These key masks will be sent to the video processor $\mathcal{F}_{vid}$, i.e. SAM2 [60], to predict the per-instance mask sequences. We formulate this process as:
@@
Figure 4: ThinkVideo for Online Reasoning Video Object Segmentation.
3.3 ThinkVideo: Online Reasoning Video Object Segmentation
Our approach can be extended to handle online video streams where future frames have yet to be observed. Refer to Fig. 4. The goal of Online Reasoning Video Object Segmentation (Online Reasoning VOS) is to generate a binary mask sequence $\hat{y}=\mathcal{M}=\{m_{t}\in\mathbb{R}^{H\times W}\}_{t=1}^{T}$ from an online input video stream $\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{T}$, and text query $q$, while the model receives video frames in temporal order and makes predictions only based on frames that have already been seen. Specifically, this process can be modified from Eq. 1 as:
@@ 3.3 ThinkVideo: Online Reasoning Video Object Segmentation
The ThinkVideo framework in Online ReasoningVOS version shares a similar architecture to the original ReasoningVIS, while “asking” the keyframe selector $\mathcal{F}_{key}$ for periodic updates of the key mask. Specifically, we introduce a hyper-parameter $\xi$ for the temporal frequency of keyframe selection. We denote $I^{key}_{t}$ as the selected keyframe at time $t$, and $p_{t}$ as its corresponding frame index, respectively. When given a complex text query $q$ and an online video frame $I_{t}$, where $t=n\xi+1$ for some $n\in\mathbb{Z}$, the model sends them to $\mathcal{F}_{key}$ to justify the the keyframe selectivity $S_{t}\in\{0,1\}$, which indicates whether $I_{t}$ should be chosen as a keyframe. This process can be formulated as:
@@
for $t=1,\xi+1,2\xi+1,...$. The selected keyframe $I^{key}_{t}$ is then used to generate a mask control and guide the video processor for $\xi$ consecutive frames. We formulate this process as the following: $m_{t}=\mathbf{0}\mathrm{~{}if~{}}I^{key}_{t^{\prime}}=\tt{Null};$ otherwise,
@@
Figure 5: Qualitative results. By thinking with CoT, our ThinkVideo demonstrates the temporal reasoning ability previous works fail to achieve. We compare ThinkVideo with Reasoning VOS method VideoLISA [4] and CoT Reasoning Image Segmentation methods Seg-Zero [41] and ThinkFirst [31]. Ground-truth masks are shown in the last column for reference. Zoom in for visual details.
@@ 4 Experiments
In this section, we will empirically demonstrate the effectiveness of our proposed ThinkVideo. We begin by introducing the experimental settings in Sec. 4.1, followed by showing the qualitative and quantitative results and comparation of ThinkVideo with previous works in Sec. 4.2. Lastly, we conduct ablation study to analyze the keyframe candidates for GPT-4o and CoT thinking process in Sec. 4.3.
@@ 4.1 Experimental Settings
Implementation details. We implement our MLLM keyframe selector $\mathcal{F}_{key}$ with GPT-4o [29] (version “2024-05-13”) for the original ThinkVideo and LLaVA1.5-7B [40] for the online ThinkVideo. We also conduct experiments with the open-sourced Gemma3-12B [65] in ThinkVideo for users with limited access to the OpenAI service. We adopt Seg-Zero [41] as the reasoning image segmentation model $\mathcal{F}_{seg}$ and SAM2 [60] as the video processor $\mathcal{F}_{vid}$. We use the pre-trained weights of each module from their official implementation. We set $\xi=\lfloor\frac{T-1}{8}\rfloor+1$ for our experiments (i.e., the integer value that generates at most and closest to 8 keyframe candidates), and $\xi=4$ for online experiments. GPT-4o’s temperature is set to $0.5$, and the maximum number of output tokens is $2500$. To evaluate the VOS task with the VIS version of ThinkVideo, we take the union of all the predicted instance-level mask(s) as the final output.
Datasets. We compare ThinkVideo with previous works on multiple benchmarks, including referring VOS datasets (MeViS [20] and Refer-DAVIS-2017 [58]) and reasoning VOS datasets (ReVOS [82] and ReasonVOS [4]). To evaluate reasoning VOS with temporally sensitive queries, we manually sample a subset of ReasonVOS, named as T-ReasonVOS, which contains only difficult examples involving time-sensitive queries. Refer to Appendix D for more dataset descriptions.
Evaluation metrics. We follow previous works and evaluate the ThinkVideo performance with region similarity ($\mathcal{J}$), contour accuracy ($\mathcal{F}$), and their average value ($\mathcal{J\&F}$). Specifically, $\mathcal{J}$ evaluates how well the pixels of the predicted mask match the ground truth overall using intersection over union (IoU), and $\mathcal{F}$ combines the precision and recall of boundary pixels, emphasizing the accuracy at the object boundaries.
Table 1: Referring VOS results on MeViS benchmark [20].
[…]
@@
URVOS [62] 27.8 25.7 29.9
LBDT [23] 29.3 27.8 30.8
MTTR [9] 30.0 28.8 31.2
ReferFormer [77] 31.0 29.8 32.2
VLT+TC [22] 35.5 33.6 37.3
LMPM [20] 37.2 34.2 40.2
@@
LISA-7B [34] 37.2 35.1 39.4
LISA-13B [34] 37.9 35.8 40.0
TrackGPT-7B [94] 40.1 37.6 42.6
TrackGPT-13B [94] 41.2 39.2 43.1
VideoLISA (OTSA) [4] 42.3 39.4 45.2
VideoLISA (Po) [4] 44.4 41.3 47.6
VISA-7B [82] 43.5 40.7 46.3
VISA-13B [82] 44.5 41.8 47.1
ThinkVideo-LLaVA1.5-7B 52.5 49.1 55.9
ThinkVideo-Gemma3-12B 52.6 48.2 57.0
ThinkVideo-GPT-4o 60.1 56.0 64.2
Table 2: Referring VOS results on Refer-DAVIS-17 benchmark [58].
[…]
@@
URVOS [62] 51.6 47.2 55.9
YOFO [35] 53.3 48.8 57.8
MLSA [74] 57.9 53.8 62.0
ReferFormer [77] 61.1 58.1 64.1
SgMg [49] 63.3 60.6 66.0
OnlineRefer [75] 64.8 61.6 67.7
@@
LISA-7B [34] 64.8 62.2 67.3
LISA-13B [34] 66.0 63.2 68.8
TrackGPT-7B [94] 63.2 59.4 67.0
TrackGPT-13B [94] 66.5 62.7 70.4
VideoLISA (OTSA) [4] 67.7 63.8 71.5
VideoLISA (Po) [4] 68.8 64.9 72.7
VISA-7B [82] 69.4 66.3 72.5
VISA-13B [82] 70.4 67.0 73.8
HyperSeg [71] 71.2 - -
ThinkVideo-LLaVA1.5-7B 73.9 70.4 77.5
ThinkVideo-Gemma3-12B 74.6 70.9 78.3
ThinkVideo-GPT-4o 79.1 75.3 82.9
Table 3: Reasoning VOS results on ReasonVOS benchmark [4]. OTSA means One-Token-Seg-All; Po means post-optimization.
[…]
Traditional methods without reasoning ability
MTTR [9] 31.1 29.1 33.1
ReferFormer [77] 32.9 30.2 35.6
SOC [44] 35.9 33.3 38.5
SgMg [49] 36.2 33.7 38.7
OnlineRefer [75] 38.7 34.6 42.9
LLM-based methods with reasoning ability
LISA [34] 31.1 29.1 33.1
VideoLISA (OTSA) [4] 45.1 43.1 47.1
VideoLISA (Po) [4] 47.5 45.1 49.9
ThinkVideo-LLaVA1.5-7B 52.0 49.5 54.5
ThinkVideo-Gemma3-12B 50.7 47.5 54.0
ThinkVideo-GPT-4o 65.5 62.4 68.7
Table 4: Reasoning VOS results on ReVOS benchmark [82]. (IT) means instruction tuning with the ReVOS training set.
[…]
Traditional methods without reasoning ability
MTTR [9] 25.5 25.1 25.9
LMPM [20] 26.4 21.2 31.7
ReferFormer [77] 28.1 26.2 29.9
LLaMA-VID [38]+LMPM 26.1 20.9 31.4
LLM-based methods with reasoning ability
LISA-13B [34] 41.6 39.8 43.5
TrackGPT-7B (IT) [94] 43.6 41.8 45.5
TrackGPT-13B (IT) [94] 45.0 43.2 46.8
VISA [82] 47.5 45.3 49.7
VISA (IT) [82] 50.9 48.8 52.9
HyperSeg [71] 55.7 53.1 58.4
ThinkVideo-LLaVA1.5-7B 46.2 43.5 48.8
ThinkVideo-Gemma3-12B 47.1 43.4 50.9
ThinkVideo-GPT-4o 55.9 52.8 59.0
4.2 Main Results
Referring VOS. We report the quantitative comparisons of ThinkVideo with previous works on the tasks of Referring VOS in Tab. 2 and Tab. 2. Online methods are highlighted in gray in these tables. Results show that both of the online and offline versions of ThinkVideo significantly outperform the state of the art. ThinkVideo achieves up to $11.3\%$ improvements over the online approaches and $15.6\%$ over the offline approaches. Notably, online ThinkVideo surpasses offline baselines in most cases, while given a more challenging task setting.
@@ 4.2 Main Results
[…]
LISA [34] 31.0 28.6 33.3
ThinkFirst [31]+SAM2 33.6 31.1 36.1
VideoLISA (OTSA) [4] 37.4 35.1 39.6
ThinkVideo-Gemma3-12B 50.2 46.8 53.6
ThinkVideo-GPT-4o 55.5 51.2 59.9
@@
Table 7: Ablation study for online keyframe selector on Refer-DAVIS-17. * varies depending on the OpenAI server busyness. is implemented on an RTX4090.
Model $\mathcal{J\&F}$ Avg. Time (sec.)
Qwen-2.5-VL-3B [3] 72.5 0.18
LLaVA1.5-7B [40] 73.9 0.24
GPT-4o [29] 73.2 0.27*
Reasoning VOS. We demonstrate the quantitative comparison of the ReasonVOS benchmark and ReVOS benchmark in Tab. 4 and Tab. 4, respectively, where ThinkVideo achieves up to $18\%$ than state-of-the-art offline methods and up to $13.3\%$ than state-of-the-art online methods, showcasing the better reasoning capability of “thinking” first. Implementation based on Gemma3-12B performs slightly worse results than in GPT-4o, possibly due to the capability of temporal reasoning and image understanding. We further report the qualitative performance of reasoning VOS with temporal-related queries in Fig. 5. When the temporal reasoning capability is needed, ThinkVideo achieves significantly better performance than any existing works. This is also verified in the quantitative results of T-ReasonVOS dataset, as shown in Tab. 7.
Reasoning VIS and online Reasoning VOS. We demonstrate the extended features of ThinkVideo in Fig. 6. In specific, the original ThinkVideo is able to generate instance-level mask sequences. The online version of ThinkVideo can update the object of interest during the tracking process. These extensions provide ThinkVideo with more flexibility to broader audience.

### v4

@@ 3.2 CoT-RVS: Reasoning Video Instance Segmentation
Following by the extraction of temporal-semantic correlation within a video, where time sensitivity has been resolved, we can now incorporate the image-based segmentation model and video processor to generate temporal-consistent mask sequence. In particular, we use the reasoning image segmentation model $\mathcal{F}_{seg}$ to generate per-instance key masks $\{\tilde{m}_{i}\in\mathbb{R}^{H\times W}\}$ based on the the instance-level selectivity. These key masks will be sent to the video processor $\mathcal{F}_{vid}$, i.e. SAM2 (Ravi et al., 2024), to predict the per-instance mask sequences. We formulate this process as:
@@
Figure 5: CoT-RVS for Online Reasoning Video Object Segmentation.
3.3 CoT-RVS: Online Reasoning Video Object Segmentation
Our approach can be extended to handle online video streams where future frames have yet to be observed. Refer to Fig.˜5. The goal of Online Reasoning Video Object Segmentation (Online Reasoning VOS) is to generate a binary mask sequence $\hat{y}=\mathcal{M}=\{m_{t}\in\mathbb{R}^{H\times W}\}_{t=1}^{T}$ from an online input video stream $\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{T}$, and text query $q$, while the model receives video frames in temporal order and makes predictions only based on frames that have already been seen. Specifically, this process can be modified from Eq.˜1 as:
@@ 3.3 CoT-RVS: Online Reasoning Video Object Segmentation
The CoT-RVS framework in Online ReasoningVOS version shares a similar architecture to the original ReasoningVIS, while “asking” the keyframe selector $\mathcal{F}_{key}$ for periodic updates of the key mask. Specifically, we introduce a hyper-parameter $\xi$ for the temporal frequency of keyframe selection. We denote $I^{key}_{t}$ as the selected keyframe at time $t$, and $p_{t}$ as its corresponding frame index, respectively. When given a complex text query $q$ and an online video frame $I_{t}$, where $t=n\xi+1$ for some $n\in\mathbb{Z}$, the model sends them to $\mathcal{F}_{key}$ to justify the the keyframe selectivity $S_{t}\in\{0,1\}$, which indicates whether $I_{t}$ should be chosen as a keyframe. This process can be formulated as:
@@
for $t=1,\xi+1,2\xi+1,$.... Refer to Appx.˜B for the prompting details for Eq.˜9. The selected keyframe $I^{key}_{t}$ is then used to generate a mask control and guide the video processor for $\xi$ consecutive frames. We formulate this process as the following: $m_{t}=\mathbf{0}\mathrm{~if~}I^{key}_{t^{\prime}}=\tt{Null};$ otherwise,
@@
Figure 6: Qualitative results. By thinking with CoT, our CoT-RVS demonstrates the temporal reasoning ability previous works fail to achieve. We compare CoT-RVS with Reasoning VOS method VideoLISA (Bai et al., 2024) and CoT Reasoning Image Segmentation methods Seg-Zero (Liu et al., 2025) and ThinkFirst (Kao et al., 2025).
@@ 4 Experiments
In this section, we will empirically demonstrate the effectiveness of our proposed CoT-RVS. We begin by introducing the experimental settings in Sec.˜4.1, followed by showing the qualitative and quantitative results and comparation of CoT-RVS with previous works in Sec.˜4.2. Lastly, we conduct ablation study to analyze the keyframe candidates for GPT-4o and CoT thinking process in Sec.˜4.3, where more ablation study, e.g. robustness, time and latency analysis, will be provided in the appendix.
Table 1: Referring VOS results on MeViS benchmark (Ding et al., 2023a).
Methods Online $\mathcal{J\&F}$ $\mathcal{J}$ $\mathcal{F}$
Traditional methods without reasoning ability
URVOS (Seo et al., 2020) 27.8 25.7 29.9
LBDT (Ding et al., 2022) 29.3 27.8 30.8
MTTR (Botach et al., 2022b) 30.0 28.8 31.2
ReferFormer (Wu et al., 2022c) 31.0 29.8 32.2
VLT+TC (Ding et al., 2021) 35.5 33.6 37.3
LMPM (Ding et al., 2023a) 37.2 34.2 40.2
LLM-based methods with reasoning ability
LISA-7B (Lai et al., 2023) 37.2 35.1 39.4
LISA-13B (Lai et al., 2023) 37.9 35.8 40.0
TrackGPT-7B (Zhu et al., 2023) 40.1 37.6 42.6
TrackGPT-13B (Zhu et al., 2023) 41.2 39.2 43.1
VideoLISA (OTSA)* (Bai et al., 2024) 42.3 39.4 45.2
VideoLISA (Po)* (Bai et al., 2024) 44.4 41.3 47.6
VISA-7B (Yan et al., 2024) 43.5 40.7 46.3
VISA-13B (Yan et al., 2024) 44.5 41.8 47.1
SAMWISE (Cuttano et al., 2024) 49.5 46.6 52.4
GLUS (Lin et al., 2025) 51.3 48.5 54.2
CoT-RVS-LLaVA1.5-7B 45.9 42.7 49.1
CoT-RVS-Gemma3-12B 44.2 40.3 48.1
CoT-RVS-GPT-4o 52.2 48.7 55.7
* OTSA: One-Token-Seg-All; Po: post-optimization.
Table 2: Referring VOS results on Refer-DAVIS-17 benchmark (Pont-Tuset et al., 2017).
Methods Online $\mathcal{J\&F}$ $\mathcal{J}$ $\mathcal{F}$
Traditional methods without reasoning ability
URVOS (Seo et al., 2020) 51.6 47.2 55.9
YOFO (Li et al., 2022) 53.3 48.8 57.8
MLSA (Wu et al., 2022a) 57.9 53.8 62.0
ReferFormer (Wu et al., 2022c) 61.1 58.1 64.1
SgMg (Miao et al., 2023b) 63.3 60.6 66.0
OnlineRefer (Wu et al., 2023b) 64.8 61.6 67.7
LLM-based methods with reasoning ability
LISA-7B (Lai et al., 2023) 64.8 62.2 67.3
LISA-13B (Lai et al., 2023) 66.0 63.2 68.8
TrackGPT-7B (Zhu et al., 2023) 63.2 59.4 67.0
TrackGPT-13B (Zhu et al., 2023) 66.5 62.7 70.4
VideoLISA (OTSA)* (Bai et al., 2024) 67.7 63.8 71.5
VideoLISA (Po)* (Bai et al., 2024) 68.8 64.9 72.7
VISA-7B (Yan et al., 2024) 69.4 66.3 72.5
VISA-13B (Yan et al., 2024) 70.4 67.0 73.8
SAMWISE (Cuttano et al., 2024) 70.6 67.4 74.5
HyperSeg (Wei et al., 2024) 71.2 - -
CoT-RVS-LLaVA1.5-7B 73.9 70.4 77.5
CoT-RVS-Gemma3-12B 74.6 70.9 78.3
CoT-RVS-GPT-4o 79.1 75.3 82.9
* OTSA: One-Token-Seg-All; Po: post-optimization.
@@ 4.1 Experimental Settings
Implementation details. We implement our MLLM keyframe selector $\mathcal{F}_{key}$ with GPT-4o (Hurst et al., 2024) (version “2024-05-13”) for the original CoT-RVS and LLaVA1.5-7B (Liu et al., 2023a) for the online CoT-RVS. We also conduct experiments with the open-sourced Gemma3-12B (Team et al., 2025) in CoT-RVS for users with limited access to the OpenAI service. Quantitative studies for different MLLMs are provided in Appendices.˜F and G to analyze our implementation choices. We adopt Seg-Zero (Liu et al., 2025) as the reasoning image segmentation model $\mathcal{F}_{seg}$ and SAM2 (Ravi et al., 2024) as the video processor $\mathcal{F}_{vid}$, where alternatives with LISA (Lai et al., 2023) and Cutie (Cheng et al., 2024) are also demonstrated in Appx.˜H for modularity analysis. We use the pre-trained weights of each module from their official implementation. We set $\xi=\lfloor\frac{T-1}{8}\rfloor+1$ for our experiments (i.e., the integer value that generates at most and closest to 8 keyframe candidates), and $\xi=4$ for online experiments. GPT-4o’s temperature is set to $0.5$, and the maximum number of output tokens is $2500$. To evaluate the VOS task with the VIS version of CoT-RVS, we take the union of all the predicted instance-level mask(s) as the final output.
Datasets. We compare CoT-RVS with previous works on multiple benchmarks, including referring VOS datasets (MeViS (Ding et al., 2023a) and Refer-DAVIS-2017 (Pont-Tuset et al., 2017)) and reasoning VOS datasets (ReVOS (Yan et al., 2024) and ReasonVOS (Bai et al., 2024)). To evaluate reasoning VOS with temporally sensitive queries, we manually sample a subset of ReasonVOS, named as T-ReasonVOS, which contains only difficult examples involving time-sensitive queries. Refer to Appx.˜D for more dataset descriptions.
Evaluation metrics. We follow previous works and evaluate the CoT-RVS performance with region similarity ($\mathcal{J}$), contour accuracy ($\mathcal{F}$), and their average value ($\mathcal{J\&F}$). Specifically, $\mathcal{J}$ evaluates how well the pixels of the predicted mask match the ground truth overall using intersection over union (IoU), and $\mathcal{F}$ combines the precision and recall of boundary pixels, emphasizing the accuracy at the object boundaries.
Table 3: Reasoning VOS results on ReasonVOS benchmark (Bai et al., 2024).
Methods Online $\mathcal{J\&F}$ $\mathcal{J}$ $\mathcal{F}$
@@
MTTR (Botach et al., 2022b) 31.1 29.1 33.1
ReferFormer (Wu et al., 2022c) 32.9 30.2 35.6
SOC (Luo et al., 2023) 35.9 33.3 38.5
SgMg (Miao et al., 2023b) 36.2 33.7 38.7
OnlineRefer (Wu et al., 2023b) 38.7 34.6 42.9
@@
LISA (Lai et al., 2023) 31.1 29.1 33.1
VideoLISA (OTSA)* (Bai et al., 2024) 45.1 43.1 47.1
VideoLISA (Po)* (Bai et al., 2024) 47.5 45.1 49.9
GLUS (Lin et al., 2025) 49.9 47.5 52.4
CoT-RVS-LLaVA1.5-7B 52.0 49.5 54.5
CoT-RVS-Gemma3-12B 50.7 47.5 54.0
CoT-RVS-GPT-4o 65.5 62.4 68.7
* OTSA: One-Token-Seg-All; Po: post-optimization.
Table 4: Reasoning VOS results on ReVOS benchmark (Yan et al., 2024).
Methods Online $\mathcal{J\&F}$ $\mathcal{J}$ $\mathcal{F}$
@@
MTTR (Botach et al., 2022b) 25.5 25.1 25.9
LMPM (Ding et al., 2023a) 26.4 21.2 31.7
ReferFormer (Wu et al., 2022c) 28.1 26.2 29.9
LLaMA-VID (Li et al., 2023c)+LMPM 26.1 20.9 31.4
@@
LISA-13B (Lai et al., 2023) 41.6 39.8 43.5
TrackGPT-7B (IT)* (Zhu et al., 2023) 43.6 41.8 45.5
TrackGPT-13B (IT)* (Zhu et al., 2023) 45.0 43.2 46.8
VISA (Yan et al., 2024) 47.5 45.3 49.7
VISA (IT)* (Yan et al., 2024) 50.9 48.8 52.9
GLUS (Lin et al., 2025) 54.9 52.4 57.3
HyperSeg (Wei et al., 2024) 55.7 53.1 58.4
CoT-RVS-LLaVA1.5-7B 46.2 43.5 48.8
CoT-RVS-Gemma3-12B 47.1 43.4 50.9
CoT-RVS-GPT-4o 55.9 52.8 59.0
* (IT): instruction tuning with the ReVOS training set.
Figure 7: Visual results of the original Reasoning VIS and its online version for Reasoning VOS. The original CoT-RVS can be used for instance-level reasoning video segmentation, while the online extension is preferred when the object of interest changes during the tracking process. Zoom in for details and refer to the appendix for more examples.
@@
Experiments $\mathcal{J\&F}$ $\mathcal{J}$ $\mathcal{F}$
LISA 31.0 28.6 33.3
ThinkFirst+SAM2 33.6 31.1 36.1
VideoLISA (OTSA) 37.4 35.1 39.6
CoT-RVS-Gemma3-12B 50.2 46.8 53.6
CoT-RVS-GPT-4o 55.5 51.2 59.9
@@
4.2 Main Results
Referring VOS. We report the quantitative comparisons of CoT-RVS with previous works on the tasks of Referring VOS in Tab.˜2 and Tab.˜2. Online methods are highlighted in gray in these tables. Results show that both of the online and offline versions of CoT-RVS significantly outperform the state of the art. CoT-RVS achieves up to $3.5\%$ improvements over the online approaches and $15.6\%$ over the offline approaches. Notably, online CoT-RVS surpasses offline baselines in most cases, while given a more challenging task setting.
Reasoning VOS. We demonstrate the quantitative comparison of the ReasonVOS benchmark and ReVOS benchmark in Tab.˜4 and Tab.˜4, respectively, where CoT-RVS achieves up to $18\%$ than state-of-the-art offline methods and up to $13.3\%$ than state-of-the-art online methods, showcasing the better reasoning capability of “thinking” first. Implementation based on Gemma3-12B performs slightly worse results than in GPT-4o, possibly due to the capability of temporal reasoning and image understanding. We further report the qualitative performance of reasoning VOS with temporal-related queries in Fig.˜6. When the temporal reasoning capability is needed, CoT-RVS achieves significantly better performance than any existing works. This is also verified in the quantitative results of T-ReasonVOS dataset, as shown in Tab.˜6.
Reasoning VIS and online Reasoning VOS. We demonstrate the extended features of CoT-RVS in Fig.˜7. In specific, the original CoT-RVS is able to generate instance-level mask sequences. The online version of CoT-RVS can update the object of interest during the tracking process. These extensions provide CoT-RVS with more flexibility to broader audience.